# Stochastic Semi-Gradient Descent for Learning Mean Field Games with Population-Aware Function Approximation

### v1

@@ 3 Linear mean field games
From this point, we remove the assumption that the state-action space is finite. In Section 2, we have shown that it is beneficial to treat the value function and population measure as a unified parameter governing the MDP dynamics. As a result, linear MDP [19] and LFA for the value function can be naturally extended to address population measure on continuous state spaces, which is crucial in various applications such as autonomous driving [9], flocking [31] and crowd modelling [20, 7]. We now introduce linear MFGs, a class of MFGs with an underlying linear MDP and a linear structure on the transition kernel w.r.t. the population measure.
@@
An MDP $(\mathcal{S},\mathcal{A},P,r,\gamma)$ is a linear MDP [19] with feature map $\phi\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d_{1}}$ if there exists $d_{1}$ (signed) population-dependent measures ${\omega}_{M}=(\omega_{M}^{(i)})_{i=1}^{d_{1}}\in\mathcal{M}(\mathcal{S})^{d_{1}}$ and an unknown population-dependent vector ${\nu}_{M}\in\mathbb{R}^{d_{1}}$, such that for any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ and population distribution $M\in\mathcal{D}(\mathcal{S})$, we have
@@
A linear MFG assumes the MDP components, the transition kernel and reward function, lie in a linear function space with known basis functions. This indicates the linear structure of the value function and population measure.
@@
The proof of Proposition 1 is deferred to Appendix C. Remarkably, requiring that the transition kernel being linear w.r.t. a measure basis is essential for the linear structure of the population measure. In this paper, we reserve letters $\theta$ for the value function parameter and $\eta$ for the population measure parameter. And we denote the concatenated parameter as $\xi=(\eta;\theta)$. Additionally, we will use parameters ($\eta$, $\theta$, $\xi$) and their corresponding functions ($Q$, $M$, $(Q,M)$) interchangeably in our analysis. For example, we say $\xi\in\mathbb{R}^{d_{1}+d_{2}}$ is a mean field equilibrium (MFE) if its corresponding value function $Q=\left<\phi,\theta\right>$ and popumeasure $M=\left<\psi,\eta\right>$ satisfy Definition 1.
@@
Example 2 implies that all the analysis for linear MFGs applies to finite MFGs (see also Appendix H).
@@ 4 SemiSGD with population-aware linear function approximation
Section 3 introduce the first linear parameterization of the population measure in MFGs, necessitating tailored stochastic update rule for the population measure estimate. We now derive the semi-gradient for the population measure with LFA. Let $M_{*}$ be the objective population measure and define the loss function $\mathcal{L}\coloneqq\frac{1}{2}\|M-M_{*}\|^{2}_{L_{2}}$. Then, the gradient of the loss is
@@
The simplex projection has a worst-case complexity of $O(d_{2}^{2})$ [11]. In semi-gradient evaluation, we note that ${G}_{\psi}$ is a constant matrix, and thus we only need to evaluate $\psi$ at the next state $s^{\prime}$ and multiply parameter with a fixed matrix, with the matrix-vector multiplication having a complexity of $O(d_{2}^{2})$. Therefore, the total worst-case operation complexity of the update rule Equation 11 is $O(d_{2}^{2})$. Additionally, the simplex projection has an expected complexity of $O(d_{2})$ [11]. And if ${G}_{\psi}$ has precomputed properties and/or special structures, such as being sparse with $O(d_{2})$ non-zero elements, the complexity of the semi-gradient evaluation is $O(d_{2})$, giving a total expected complexity of $O(d_{2})$, which is the same as the operation complexity of Equation 3 for finite MFGs with a state space of size $|\mathcal{S}|=O(d_{2})$. In summary, Equation 11 is as simple and efficient as the population measure update for finite MFGs.
Similarly, we can get the TD update rule for the value function parameter $\theta$[32]:
@@
where we denote $\mathfrak{g}_{t}(\theta)\coloneqq\mathfrak{g}(\theta;O_{t})$, $O_{t}=(s_{t},a_{t},r_{t},s_{t+1},a_{t+1})$, and $\Pi_{D}$ is the projection operator onto the Euclidean ball $B_{D}^{d_{1}}\coloneqq\{\theta\in\mathbb{R}^{d_{1}}:\|\theta\|\leq D\}$. Similar to the population measure update, the projection is commonly used with LFA to ensure that the value function parameter remains in a bounded region [5, 40], which is automatically satisfied in finite MFGs Appendix H. Specifically, we need $D\geq\|\theta_{*}\|$, where $\theta_{*}$ is the MFE value function parameter.
@@
$\xi$ is an MFE if and only if $\bar{\mathfrak{g}}_{\xi}(\xi)=0$.
We prove an extended version of Proposition 2 in Appendix D, showing that the stationary point actually corresponds to a projected MFE for general (non-linear) MFGs.
@@ 5 Sample complexity analysis
Assumptions 1 and 2 are regularity conditions ensuring learning dynamics’ smoothness, i.e., the updates are not too drastic between iterations. Assumption 3 is a standard assumption in for online methods with Markovian sampling [5, 40, 2, 36], ensuring the environment is relatively easy to explore. We can now bound the mean squared error of SemiSGD recursively, which directly gives several finite-time error bounds.
@@
Let $\xi_{*}=(\theta_{*};\eta_{*})$ be an MFE parameter. Let $\{\xi_{t}=(\theta_{t};\eta_{t})\}$ be a sequence of parameters generated by SemiSGD. Then, under Assumptions 1, 2 and 3, there exists $w\in(0,1/2]$ such that if the Lipschitz constants are small enough ($3\sigma\max\{L_{P},L_{\pi}\}H+L_{r}\leq 2w$, see Lemma 9 in Section F.1), we have
$\mathbb{E}\!\left\|\xi_{t+1}\!\!-\!\xi_{*}\right\|^{2}\!\!\leq\!\!(1-\alpha_{t}w)\mathbb{E}\|\xi_{t}-\xi_{*}\|^{2}+H^{2}\cdot O(\alpha_{t}^{2}\log\alpha^{-1}_{t})+\frac{\max\{L_{P},\!L_{\pi},\!L_{r}\}^{2}\!H^{4}}{w}\cdot O(\alpha_{t}^{3}\log^{4}\alpha_{t}^{-1}),$ (16)
where $H\coloneqq(1+\gamma)D+R+2F$ can be seen as the scale of the problem.
@@
Let $\alpha_{0}=(\log T)/(wT)$. Then Corollary 1 states $\mathbb{E}\|\xi_{T}-\xi_{*}\|^{2}=O\left((H^{2}\log^{2}T)/(w^{2}T)\right)$, giving an $O(\epsilon^{-2}\log^{2}\epsilon^{-1})$ sample complexity for an $\epsilon$-MFE ($\mathbb{E}\|\xi_{T}-\xi_{*}\|\leq\epsilon$). A linearly decaying step-size sequence improves the logarithmic factor, giving a convergence rate of $O\left((H^{2}\log T)/(w^{2}T)\right)$ (see Corollary 3). Reducing to finite MFGs, the complexity becomes $\widetilde{O}\left(\frac{SAR^{2}}{\lambda^{2}(1-\gamma)^{4}T}\right)$ (see Corollary 4).
Remark 2 (Small Lipschitz constants).
SemiSGD follows a non-stationary Markov chain with its kernel determined by the value function and population measure, which are updated on the go. When the kernel is sensitive to the parameter update (large Lipschitz constants), the Markov chain is rapidly changing [39], whose convergence is not well understood. As a result, for theoretical analysis, we require the Lipschitz constants to be small (see Lemma 9), so the descent direction (attraction property of MFE) dominants the non-stationarity. Such requirements are not implemented in the experiments.
@@ 6 Approximation error for non-linear MFGs
Let $\xi_{\diamond}=(\theta_{\diamond};\eta_{\diamond})$ be the convergence point of SemiSGD with PA-LFA (i.e., the stationary point w.r.t. the semi-gradient, or the projected MFE). Let $(q_{*},\mu_{*})$ be the actual MFE. We write $\mathcal{P}_{\diamond}\coloneqq\mathcal{P}_{\xi_{\diamond}}$ for short, and similarly for $\mathcal{P}_{*}$, $\mathcal{T}_{\diamond}$, and $\mathcal{T}_{*}$. We have the following approximation error characterization.
@@
Assumption 3 indicates the existence of $k\in\mathbb{N}$ such that
@@
where $\epsilon_{\phi}\coloneqq\|q_{*}-\Pi_{\phi}q_{*}\|_{\infty}$ and $\epsilon_{\psi}\coloneqq\|\mathcal{P}_{*}-\Pi_{\psi}\mathcal{P}_{*}\|_{\mathrm{TV}}$ are inherent approximation error induced by the projection onto the linear span of basis $\phi$ and $\psi$, which is independent of the algorithm. Moreover, if $\|\mathcal{P}_{*}-\mathcal{P}_{*}^{\infty}\|_{\mathrm{TV}}=\rho$ (e.g., $\mathcal{P}_{*}$ induces a reversible Markov chain), $k=1$.
The approximation error in Theorem 2 scales with the inherent approximation error determined by the chosen basis, which is consistent with the approximation error of LFA for value function [33].
@@ 7 Numerical experiments
We conduct experiments on two MFG examples in autonomous driving scenarios [37]: speed control on a ring road and routing game on a network. We equip FPI with entropy regularization (ER), fictitious play (FP), and online mirror descent (OMD). The following parameters are shared for all methods and two experiments: total number of samples $T=10^{5}$, with the number of outer loops being $K=200$ for FPI-type methods, constant step-size $\alpha=10^{-3}$, policy operator $\Gamma_{\pi}=\operatorname{softmax}$ with a large inverse temperature (near-greedy). The convergence performance include exploitability and mean squared error (MSE). We compare distributions with the reference MFE distribution, which is calculated by model-based FPI with FP. All results are averaged over $10$ independent runs with random initialization. $95\%$ confidence regions are plotted. Please refer to Appendix A for the full setup of the experiments. In the main body, we present only the results of the speed control problem due to page limit. The details of the other game (routing) can be found in Appendix A.
Speed control on a ring road.
We consider a speed control game on a ring road $\mathcal{S}\cong\mathbb{S}^{1}$. The cost function $r(s,a,\mu)=-\left(b(s)+\left(1-\mu(s)/\mu_{\mathrm{jam}}\right)/2-a/a_{\mathrm{max}}\right)^{2}\Delta s/2$, is a variant of the Lighthill-Whitham-Richards function and a bonus term $b$ that encodes the location preference, $\mu_{\mathrm{jam}}$ is the jam density, and $a_{\mathrm{max}}$ is the maximum speed. ER is implemented with a low inverse temperature of $10^{4}$, while all other methods use an inverse temperature of $10^{9}$, guaranteeing the policies are near-greedy. The MSE, exploitabity [23], and learned population comparison are reported in Figure 1(a), Figure 2, and Figure 3, respectively. The MSE comparison demonstrates that SemiSGD 1) is more efficient, with a much faster convergence, 2) is more stable without any stabilization techniques, 3) and is more accurate, with a lower MSE. For the exploitability, SemiSGD seems to oscillate and be outperformed by ER. However, exploitability is an operator determined by the underlying MDP, which may not be smooth and does not directly reflecting the learning dynamics. The learned population distribution shows that FPI with ER is still far from the reference MFE. We also conduct a comparison between LFA and discretization, which shows that LFA is more accurate than discretization for small dimensions (Table 2; please refer to Appendix A for the full report).
@@
8 Conclusion and discussion
The proposed SemiSGD makes learning MFGs more accessible and practical in real-world applications, including but not limited to online autonomous vehicle driving, real-time pedestrian or crowd dynamics, internet packet routing, social opinion dynamics, and epidemiology. Our methodology is generalizable and can be extended to : (finite) population games, graphon games, policy optimization for MFGs, and evolutionary MDPs.
An open challenge is to relax or remove the small Lipschitz constants assumption in our theoretical analysis. As demonstrated in the experiments, SemiSGD is stable without requiring the Lipschitz constants to be small. Bridging this gap requires a deeper understanding of the rapidly changing Markov chains (RCMCs) induced by SemiSGD. It is also interesting to establish similar theoretical guarantees for monotone MFGs, a parallel condition to the smoothness assumption.

### v2

@@ 3 Linear mean field games
We now extend our setup beyond finite state-action spaces. Viewing the value function and population measure as a unified parameter, we can naturally extend linear MDP (Jin et al., 2020) and LFA for the value function to handle population measures on large or continuous state spaces. Before extending LFA to population measures in the next section, we first introduce linear MFGs, a class of MFGs based on linear MDPs with a linear structure in the transition kernel w.r.t. the population measure.
@@
An MDP $(\mathcal{S},\mathcal{A},P,r,\gamma)$ is a linear MDP (Jin et al., 2020) with feature map $\phi\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d_{1}}$ if there exists $d_{1}$ (signed) population-dependent measures ${\omega}_{M}=(\omega_{M}^{(i)})_{i=1}^{d_{1}}\in\mathcal{M}(\mathcal{S})^{d_{1}}$ and an unknown population-dependent vector ${\nu}_{M}\in\mathbb{R}^{d_{1}}$, such that for any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ and population distribution $M\in\mathcal{D}(\mathcal{S})$, we have
@@
A linear MFG assumes MDP components, the transition kernel and reward function, lie in the linear span of known basis functions, giving a linear structure to the value function and population measure.
@@
The proof of Proposition 1 is deferred to Appendix E. Remarkably, requiring that the transition kernel being linear w.r.t. a measure basis is essential for the linear structure of the population measure. In this paper, we reserve letters $\theta$ for the value function parameter and $\eta$ for the population measure parameter. And we denote the concatenated parameter as $\xi=(\eta;\theta)$. Additionally, we will use parameters ($\eta$, $\theta$, $\xi$) and their corresponding functions ($Q$, $M$, $(Q,M)$) interchangeably in our analysis. For example, we say $\xi\in\mathbb{R}^{d_{1}+d_{2}}$ is a mean field equilibrium (MFE) if its corresponding value function $Q=\left<\phi,\theta\right>$ and population measure $M=\left<\psi,\eta\right>$ satisfy Definition 1.
@@
Example 2 implies that all the analysis for linear MFGs applies to finite MFGs (see also Appendix K).
@@ 4 SemiSGD with population-aware linear function approximation
Proposition 1 presents the first linear parameterization of the population measure in MFGs. Applying this to general MFGs leads to population-aware linear function approximation (PA-LFA). More discussions on motivations of PA-LFA can be found in Section B.2. PA-LFA necessitates a tailored stochastic update rule for the population measure estimate. Let $M_{*}$ be the objective population measure and define the loss function $\mathcal{L}\coloneqq\frac{1}{2}\|M-M_{*}\|^{2}_{L_{2}}$. Then, the gradient of the loss is
@@
The simplex projection has a worst-case complexity of $O(d_{2}^{2})$ (Condat, 2016). In semi-gradient evaluation, ${G}_{\psi}$ is a constant matrix, and thus we only need to evaluate $\psi$ at the next state $s^{\prime}$ and multiply parameter with a fixed matrix, with the matrix-vector multiplication having a complexity of $O(d_{2}^{2})$. Therefore, the total worst-case operation complexity of the update rule Equation 11 is $O(d_{2}^{2})$. Moreover, the simplex projection has an expected complexity of $O(d_{2})$ (Condat, 2016). And if ${G}_{\psi}$ has precomputed properties and/or special structures, such as being sparse with $O(d_{2})$ non-zero elements, the complexity of the semi-gradient evaluation is $O(d_{2})$, giving a total expected operation complexity of $O(d_{2})$, which is the same as that of Equation 3 for finite MFGs with a state space of size $|\mathcal{S}|=O(d_{2})$.
Similarly, we can get the TD update rule for the value function parameter $\theta$ (Sutton & Barto, 2018):
@@
where we denote $\mathfrak{g}_{t}(\theta)\coloneqq\mathfrak{g}(\theta;O_{t})$, $O_{t}=(s_{t},a_{t},r_{t},s_{t+1},a_{t+1})$, and $\Pi_{D}$ is the projection operator onto the Euclidean ball $B_{D}^{d_{1}}\coloneqq\{\theta\in\mathbb{R}^{d_{1}}:\|\theta\|\leq D\}$. Similar to the population measure update, the projection is commonly used with LFA to ensure that the value function parameter remains bounded (Bhandari et al., 2018; Zou et al., 2019), which is automatically satisfied in finite MFGs (Appendix K). Specifically, we need $D\geq\|\theta_{*}\|$, where $\theta_{*}$ is the MFE value function parameter.
@@
For linear MFGs, $\xi$ is an MFE if and only if $\bar{\mathfrak{g}}_{\xi}(\xi)=0$.
We prove an extended version of Proposition 2 in Appendix F, showing that the stationary point actually corresponds to a projected MFE for general (non-linear) MFGs.
@@ 5 Sample complexity analysis
For notational convenience, we define an ergodicity constant $\sigma\coloneqq 2+\hat{n}+m\rho^{\hat{n}}/(1-\rho)$ with $\hat{n}\coloneqq\left\lceil\log_{\rho}m^{-1}\right\rceil$, and $H\coloneqq(1+\gamma)D+R+2F$, which can be regarded as the scale of the problem.
Assumption 4.
The Lipschitz constants are sufficiently small such that $3\sigma\max\{L_{P},L_{\pi}\}H+L_{r}\leq 2w$, where $w\in(0,1/2]$ is a problem-dependent constant (defined in Lemma 9 in Section H.1)
Remark 2.
Assumption 1 is a standard regularity condition for MFGs that do not assume a blanket contractivity assumption (Cui & Koeppl, 2021; Angiuli et al., 2023; Anahtarci et al., 2023; Yardim et al., 2023; Huang et al., 2024). Assumption 2 ensures the smoothness of policy updates, a condition typically met through regularization. Example policy operators satisfying Assumption 2 include softmax (Cui & Koeppl, 2021; Zaman et al., 2020; Angiuli et al., 2023), mirror descent (Perolat et al., 2021; Laurière et al., 2022b), and mirror ascent (Yardim et al., 2023). Assumption 3 is a standard mixing assumption for online methods with Markovian sampling (Bhandari et al., 2018; Zou et al., 2019; Angiuli et al., 2022; Zaman et al., 2023), ensuring that the agent’s exploration adequately covers the state-action space. Notably, Assumption 3 is also typically satisfied by regularizing behavior policies (Yardim et al., 2023). It is noteworthy that even earlier works using a blanket contractivity assumption recognize that contractivity can be achieved through regularization, provided that “the transition kernel and reward function are Lipschitz and the corresponding Lipschitz constants are small enough” (Xie et al., 2021). Therefore, Assumption 4 is the key assumption that guarantees the contractivity of learning dynamics. In summary, our assumptions closely align with those in Angiuli et al. (2023); Anahtarci et al. (2023), and are not more restrictive than those in the literature for contractive MFGs. See also Appendix I for convergence results in the absence of Assumption 4.
We can now bound the mean squared error of SemiSGD recursively, which directly gives several finite-time error bounds.
@@
Let $\xi_{*}=(\theta_{*};\eta_{*})$ be an MFE parameter. Let $\{\xi_{t}=(\theta_{t};\eta_{t})\}$ be a sequence of parameters generated by SemiSGD. Then, under Assumptions 1, 2, 3 and 4, we have
$\mathbb{E}\!\left\|\xi_{t+1}\!\!-\!\xi_{*}\right\|^{2}\!\!\leq\!\!(1-\alpha_{t}w)\mathbb{E}\|\xi_{t}-\xi_{*}\|^{2}+H^{2}\cdot O(\alpha_{t}^{2}\log\alpha^{-1}_{t})+\frac{\max\{L_{P},\!L_{\pi},\!L_{r}\}^{2}\!H^{4}}{w}\cdot O(\alpha_{t}^{3}\log^{4}\alpha_{t}^{-1}).$ (16)
@@
Let $\alpha_{0}=(\log T)/(wT)$. Then Corollary 1 states $\mathbb{E}\|\xi_{T}-\xi_{*}\|^{2}=O\left((H^{2}\log^{2}T)/(w^{2}T)\right)$, giving an $O(\epsilon^{-2}\log^{2}\epsilon^{-1})$ sample complexity for an $\epsilon$-MFE ($\mathbb{E}\|\xi_{T}-\xi_{*}\|\leq\epsilon$). A linearly decaying step-size sequence improves the logarithmic factor, giving a convergence rate of $O\left((H^{2}\log T)/(w^{2}T)\right)$ (see Corollary 3). Reducing to finite MFGs, the complexity becomes $\widetilde{O}\left(\frac{SAR^{2}}{\lambda^{2}(1-\gamma)^{4}T}\right)$ (see Corollary 4). These results also imply the uniqueness of the MFE under the assumptions of Theorem 1. The sample complexity is tight up to a logarithmic factor, consistent with other stochastic approximation methods, and strictly better than existing results for learning MFGs (Table 2). See Section B.3 for more discussions.
@@ 6 Approximation error for non-linear MFGs
Let $\xi_{\diamond}=(\theta_{\diamond};\eta_{\diamond})$ be the convergence point of SemiSGD with PA-LFA. Let $(q_{*},\mu_{*})$ be the actual MFE. Write $\mathcal{P}_{\diamond}\coloneqq\mathcal{P}_{\xi_{\diamond}}$ for short, and similarly for $\mathcal{P}_{*}$, $\mathcal{T}_{\diamond}$, and $\mathcal{T}_{*}$. Then, Assumption 3 indicates the existence of $k\in\mathbb{N}$ such that
@@
where $\epsilon_{\phi}\coloneqq\|q_{*}-\Pi_{\phi}q_{*}\|_{\infty}$ and $\epsilon_{\psi}\coloneqq\|\mathcal{P}_{*}-\Pi_{\psi}\mathcal{P}_{*}\|_{\mathrm{TV}}$ are inherent approximation error induced by the projection onto the linear spans of basis $\phi$ and $\psi$, which is independent of the algorithm. Moreover, if $\|\mathcal{P}_{*}-\mathcal{P}_{*}^{\infty}\|_{\mathrm{TV}}=\rho$ (e.g., $\mathcal{P}_{*}$ induces a reversible Markov chain), $k=1$.
The proof of Theorem 2 is given in Appendix J. The approximation error of PA-LFA scales with the inherent approximation error determined by the chosen basis, which is consistent with the approximation error of LFA for value function (Tsitsiklis & Van Roy, 1997).
@@ 7 Numerical experiments
To evaluate SemiSGD with PA-LFA, we conduct six experiments on three MFG examples: speed control (Section C.2), flocking (Section C.4), and network routing (Section C.6). When comparing with FPI-type methods, alongside vanilla online FPI, we equip FPI with entropy regularization (ER), fictitious play (FP), and mirror descent (MD). Please refer to Appendix C for the detailed setup, full results and analysis, and additional experiments. Here, we present some highlights of the results.
Figure 2(a) reports MSE curves for the speed control game. This example showcases the sample efficiency, stability, and accuracy of SemiSGD, particularly compared to vanilla FPI. The flocking game is highly sensitive to the reward (Section C.5). Large regularization may obscure the reward signal, leading to near-constant policies and population distributions, while other methods except SemiSGD cannot handle non-regularized policies well. Figure 2(b) reports exploitability curves for the flocking game, showing that only SemiSGD can effectively learn the game with near-greedy policies. The goal of the network routing game is to direct the traffic from the origin to the destination. Figure 2(c) reports the population distribution learned by SemiSGD on this highly non-smooth game. Figure 3 examines how the number of inner loop iterations ($K$) affect the convergence of FPI (Section C.3): it suggests that, given the same number of total samples, reducing $K$ monotonically improves the convergence. Notably, when $K=1$, online FPI is equivalent to SemiSGD, which removes the forward-backward structure. Figures 4 and 1 compare PA-LFA with discretization (Section C.7) on learned distribution and accuracy respectively, demonstrating that with a proper choice of measure basis, PA-LFA generalizes better and achieves higher accuracy.
(a) Speed control.
(b) Flocking.
(c) Network routing.
Figure 2: Convergence performance of SemiSGD and FPI.
@@
8 Conclusion
This work proposes a simple SGD-type method for learning MFGs, eliminating the forward-backward structure in FPI-type methods and the need for complex stabilization mechanisms. Our approach leverages a novel perspective that unifies the value function and population measure as a single parameter, enabling a straightforward analysis that elegantly handle the coupling between policy and population while offering insights into the learning dynamics.
Building on this perspective, we develop the first population-aware linear function approximation for MFGs on large or continuous state-action spaces, with sample and operation efficiency guarantees, as well as approximation error characterization. Function approximation is just one example of how techniques from policy learning can be extended to population learning using this unified perspective, and we expect further advancements in this direction.
More broadly, our methodology is generalizable to other dynamical systems that can be parameterized by learnable parameters, including (finite) population games (Sandholm, 2010), evolutionary games (Hilbe et al., 2018), graphon games (Zhou et al., 2024), and policy optimization for MFGs (Zeng et al., 2025).
Acknowledgments
This work is sponsored by NSF under CAREER award number CMMI-1943998.